# MATHWELL: Generating Educational Math Word Problems at Scale

### 24 Feb 2024 v1

@@ Abstract
Math word problems are critical K-8 educational tools, but writing them is time-consuming and requires domain expertise. We suggest that language models can support K-8 math education by automatically generating problems at scale. To be educational, generated problems must be 1) solvable, 2) accurate, and 3) appropriate. Existing datasets are unlabeled for these criteria, making them ill-suited for training problem generators. We introduce MATHWELL, a Llama-2 (70B) model iteratively finetuned to generate K-8 math word problems using data from expert annotation. Using MATHWELL, we generate the largest English word problem dataset to date, containing 20,490 problems. 3,484 are scored by domain experts who find MATHWELL has a 40% higher share of problems that have executable solutions and meet all criteria than alternatives, with 74% of its problems with executable solutions being solvable, accurate, and appropriate.
@@ 1 Introduction
We use MATHWELL to generate Synthetic Grade School Math (SGSM), a dataset of 20,490 word problems. This is the largest public set of English math word problems. We acquire annotations from domain experts for 3,484 model samples during and after training to score the quality of our data, resulting in a SGSM subset with 2,093 samples with gold labels for solvability, accuracy, and appropriateness. We find that MATHWELL is effective at generating custom math word problems without context: 74% of its output problems with executable code are simultaneously solvable, accurate, and appropriate and nearly all successfully include a specified topic. These outputs are similar to human-written math QA datasets. MATHWELL has a nearly 20% higher share of outputs with executable code that meet our three evaluation criteria than alternative LLMs. Further, we find that SGSM is more age appropriate in terms of readability than other datasets. We also show that our human annotations can be used to train text classifiers to automatically score MATHWELL outputs. Our model, data, and annotations are released publicly.111https://github.com/bryanchrist/MATHWELL
@@
We release SGSM and show it is high quality through automatic evaluation. SGSM is the only existing dataset designed to train context-free word problem generators, is the largest English math grade school dataset, and contains more readable questions than existing grade school math datasets.
@@ 3.2 Generating a Dataset with MATHWELL
We next prompt MATHWELL to generate a large dataset, which we call Synthetic Grade School Math (SGSM). SGSM contains 20,490 question/answer pairs generated by MATHWELL. SGSM has two subsets: SGSM, comprised of 2,093 MaC question/answer pairs, and SGSM, comprised of 18,397 question/answer pairs that have executable code but are not labeled for solvability, accuracy, and appropriateness. SGSM is the largest English grade school math QA dataset. To support research in automatically scoring model-generated math word problems, we release all of our annotated data. To the best of our knowledge, this is the only human annotated model-generated math word problem dataset (see Appendix B for more details).
@@
Table 1 shows the advantages of SGSM over existing datasets. First, it is the largest available dataset, which may enable further training of task-specific models. Second, SGSM is the only dataset with both Program of Thought (PoT) solutions written as Python functions and questions that assess mathematical concepts appropriate for K-8 students, which we find critical for training context-free word problem generators (see Appendix A). Based on the experiments reported in Section 4, we believe SGSM is also predominately comprised of questions that are mathematically appropriate for K-8 students. Third, SGSM and its subsets have the lowest average Flesch-Kincaid grade level (FKGL) Flesch (1948); Kincaid et al. (1975) and comparable New Dale-Chall (NDC) readability score Chall and Dale (1995), so the questions may be more appropriate for those who struggle to read.

### 28 Feb 2024 v2

@@ Abstract
Math word problems are critical K-8 educational tools, but writing them is time-consuming and requires domain expertise. We suggest that language models can support K-8 math education by automatically generating problems at scale. To be educational, generated problems must be 1) solvable, 2) accurate, and 3) appropriate. Existing datasets are unlabeled for these criteria, making them ill-suited for training problem generators. We introduce MATHWELL, a Llama-2 (70B) model iteratively finetuned to generate K-8 math word problems using data from expert annotation. Using MATHWELL, we generate the largest English word problem dataset with Program of Thought (PoT) rationales to date, containing 20,490 problems. 3,484 are scored by domain experts who find MATHWELL has a 40% higher share of problems that have executable solutions and meet all criteria than alternatives, with 74% of its problems with executable solutions being solvable, accurate, and appropriate. We release our model, data, and annotations.111https://github.com/bryanchrist/MATHWELL
@@ 1 Introduction
We use MATHWELL to generate Synthetic Grade School Math (SGSM), a dataset of 20,490 word problems. This is the largest public set of English math word problems with PoT rationales. We acquire annotations from domain experts for 3,484 model samples during and after training to score the quality of our data, resulting in a SGSM subset with 2,093 samples with gold labels for solvability, accuracy, and appropriateness. We find that MATHWELL is effective at generating custom math word problems without context: 74% of its output problems with executable code are simultaneously solvable, accurate, and appropriate and nearly all successfully include a specified topic. These outputs are similar to human-written math QA datasets. MATHWELL has a nearly 20% higher share of outputs with executable code that meet our three evaluation criteria than alternative LLMs. Further, we find that SGSM has a more age appropriate reading level than other datasets. We also show that our human annotations can be used to train text classifiers to automatically score MATHWELL outputs.
@@
We release SGSM and show it is high quality through automatic evaluation. SGSM is the only existing dataset designed to train context-free word problem generators, is the largest English math grade school dataset with PoT solutions, and contains more readable questions than existing grade school math datasets.
@@ 3.2 Generating a Dataset with MATHWELL
We next prompt MATHWELL to generate a large dataset, which we call Synthetic Grade School Math (SGSM). SGSM contains 20,490 question/answer pairs generated by MATHWELL. SGSM has two subsets: SGSM, containing 2,093 MaC question/answer pairs, and SGSM, containing 18,397 question/answer pairs that have executable code but are not labeled for solvability, accuracy, and appropriateness. SGSM is the largest English grade school PoT math QA dataset. To support research in automatically scoring model-generated math word problems, we release our annotated data. To the best of our knowledge, this is the only human annotated model-generated math word problem dataset (see Appendix B for details).
@@
Table 1 shows the advantages of SGSM over existing datasets. First, it is the largest available PoT dataset, which may enable further training of task-specific models. Second, SGSM is the only dataset with both Program of Thought (PoT) solutions written as Python functions and questions that assess mathematical concepts appropriate for K-8 students, which we find critical for training context-free word problem generators (see Appendix A). Based on the experiments reported in Section 4, we believe SGSM is also predominately comprised of questions that are mathematically appropriate for K-8 students. Third, SGSM and its subsets have the lowest average Flesch-Kincaid grade level (FKGL) Flesch (1948); Kincaid et al. (1975) and comparable New Dale-Chall (NDC) readability score Chall and Dale (1995), so the questions may be more appropriate for those who struggle to read.
@@ 5 Conclusions
Acknowledgements
The authors thank Zooniverse Zooniverse for providing a free and user-friendly platform for data annotation as well as our expert volunteer annotators for providing high-quality labels and feedback on our evaluation criteria/directions.